1/6-8/7=n/42
42.(1/6-8/7)=42.(n/42)
7-48=n
n=-41

can you please explain how my teacher got 7-48=n she does not like to explain herself and expect us to automatically get it... thank you

Answer by stanbon(75887)

1/6-8/7=n/42
Multiply both sides by 42::
42.(1/6-8/7)=42.(n/42)
42*(1/6)+42(-8/7) = 42(n/42)
----------
7 - 6*8 = n
-----
7 - 48=n
-----------
n=-41
Note: You teacher is teachig you algebra.
He/she expects that you know how to do the arithmatic.
--------
Cheers,
Stan H.
===============